\def\romsup#1{{\edef\next{\the\font}$^{\next#1}$}} \def\cprime{$'$}
  \def\cftil#1{\ifmmode\setbox7\hbox{$\accent"5E#1$}\else
  \setbox7\hbox{\accent"5E#1}\penalty 10000\relax\fi\raise 1\ht7
  \hbox{\lower1.15ex\hbox to 1\wd7{\hss\accent"7E\hss}}\penalty 10000
  \hskip-1\wd7\penalty 10000\box7} \def\Dbar{\leavevmode\lower.6ex\hbox to
  0pt{\hskip-.23ex \accent"16\hss}D}
  \def\cfac#1{\ifmmode\setbox7\hbox{$\accent"5E#1$}\else
  \setbox7\hbox{\accent"5E#1}\penalty 10000\relax\fi\raise 1\ht7
  \hbox{\lower1.15ex\hbox to 1\wd7{\hss\accent"13\hss}}\penalty 10000
  \hskip-1\wd7\penalty
  10000\box7}\def\cfudot#1{\ifmmode\setbox7\hbox{$\accent"5E#1$}\else
  \setbox7\hbox{\accent"5E#1}\penalty 10000\relax\fi\raise 1\ht7
  \hbox{\raise.1ex\hbox to 1\wd7{\hss.\hss}}\penalty 10000 \hskip-1\wd7\penalty
  10000\box7} \def\polhk#1{\setbox0=\hbox{#1}{\ooalign{\hidewidth
  \lower1.5ex\hbox{`}\hidewidth\crcr\unhbox0}}}
  \def\cydot{\leavevmode\raise.4ex\hbox{.}}
  \def\cfgrv#1{\ifmmode\setbox7\hbox{$\accent"5E#1$}\else
  \setbox7\hbox{\accent"5E#1}\penalty 10000\relax\fi\raise 1\ht7
  \hbox{\lower1.05ex\hbox to 1\wd7{\hss\accent"12\hss}}\penalty 10000
  \hskip-1\wd7\penalty 10000\box7}
  \def\uarc#1{\ifmmode{\lineskiplimit=0pt\oalign{$#1$\crcr
  \hidewidth\setbox0=\hbox{\lower1ex\hbox{{\rm\char"15}}}\dp0=0pt
  \box0\hidewidth}}\else{\lineskiplimit=0pt\oalign{#1\crcr
  \hidewidth\setbox0=\hbox{\lower1ex\hbox{{\rm\char"15}}}\dp0=0pt
  \box0\hidewidth}}\relax\fi}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{AHJM88b}

\bibitem[AHJM88a]{MR935523}
J.~F. Adams, J.-P. Haeberly, S.~Jackowski, and J.~P. May, \emph{A
  generalization of the {A}tiyah-{S}egal completion theorem}, Topology
  \textbf{27} (1988), no.~1, 1--6. \MR{MR935523 (90e:55026)}

\bibitem[AHJM88b]{MR935524}
\bysame, \emph{A generalization of the {S}egal conjecture}, Topology
  \textbf{27} (1988), no.~1, 7--21. \MR{935524 (90e:55027)}

\bibitem[Ati68]{MR0228000}
M.~F. Atiyah, \emph{Bott periodicity and the index of elliptic operators},
  Quart. J. Math. Oxford Ser. (2) \textbf{19} (1968), 113--140. \MR{MR0228000
  (37 \#3584)}

\bibitem[{Eme}10]{2010arXiv1004.2970E}
H.~{Emerson}, \emph{{Localization techniques in circle-equivariant KK-theory}},
  ArXiv e-prints (2010).

\bibitem[Jac77]{MR0448377}
Stefan Jackowski, \emph{Equivariant {$K$}-theory and cyclic subgroups},
  Transformation groups (Proc. Conf., Univ. Newcastle upon Tyne, Newcastle upon
  Tyne, 1976), Cambridge Univ. Press, Cambridge, 1977, pp.~76--91. London Math.
  Soc. Lecture Note Series, No. 26. \MR{MR0448377 (56 \#6684)}

\bibitem[Kas88]{MR918241}
G.~G. Kasparov, \emph{Equivariant {$KK$}-theory and the {N}ovikov conjecture},
  Invent. Math. \textbf{91} (1988), no.~1, 147--201. \MR{MR918241 (88j:58123)}

\bibitem[May96]{MR1413302}
J.~P. May, \emph{Equivariant homotopy and cohomology theory}, CBMS Regional
  Conference Series in Mathematics, vol.~91, Published for the Conference Board
  of the Mathematical Sciences, Washington, DC, 1996, With contributions by M.
  Cole, G. Comezana, S. Costenoble, A. D. Elmendorf, J. P. C. Greenlees, L. G.
  Lewis, Jr., R. J. Piacenza, G. Triantafillou, and S. Waner. \MR{MR1413302
  (97k:55016)}

\bibitem[McC86]{MR862427}
James~E. McClure, \emph{Restriction maps in equivariant {$K$}-theory}, Topology
  \textbf{25} (1986), no.~4, 399--409. \MR{862427 (88f:55022)}

\bibitem[MM04]{MR2083579}
Michel Matthey and Guido Mislin, \emph{Equivariant {$K$}-homology and
  restriction to finite cyclic subgroups}, $K$-Theory \textbf{32} (2004),
  no.~2, 167--179. \MR{2083579 (2005k:19012)}

\bibitem[Sch92]{MR1193150}
Claude Schochet, \emph{On equivariant {K}asparov theory and
  {S}panier-{W}hitehead duality}, $K$-Theory \textbf{6} (1992), no.~4,
  363--385. \MR{MR1193150 (94c:19006)}

\bibitem[Seg68]{MR0248277}
Graeme Segal, \emph{The representation ring of a compact {L}ie group}, Inst.
  Hautes \'Etudes Sci. Publ. Math. (1968), no.~34, 113--128. \MR{MR0248277 (40
  \#1529)}

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{mathrsfs}
\usepackage{enumerate}
\usepackage[english]{babel}
\usepackage{hyperref}
\usepackage[all,cmtip]{xy}
\entrymodifiers={+!!<0pt,\fontdimen22\textfont2>}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}[thm]{Lemma}
\newtheorem*{thm-i}{Theorem}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{rem}[thm]{Remark}
\newtheorem{defn}[thm]{Definition}

\newtheorem*{acknowledgments*}{Acknowledgments}

\numberwithin{equation}{section}

\theoremstyle{remark}

\mathchardef\ordinarycolon\mathcode`\: 
\def\vcentcolon{\mathrel{\mathop\ordinarycolon}} 
\providecommand*\coloneqq{\mathrel{\vcentcolon\mkern-1.2mu}=}

\def\Z{{\mathbb Z}} 

\DeclareMathOperator\res{res}

\def\KK{{K\!K}}
\def\K{{\widetilde{K}}}
\def\E{{\widetilde E}}

\def\L{{\widetilde E}}

\def\F{{\mathcal F}}
\def\H{{\mathcal H}}

\def\I{{\mathcal I}}
\def\C{{\mathcal C}}
\def\O{{\mathcal O}}

\def\cast{$C^{*}$}
\def\pt{*}

\newcommand\thmA{(2)}
\newcommand\thmC{(1)}

\begin{document}

\title{Restriction maps in equivariant $\KK$-theory.}
\author{Otgonbayar Uuye}
\date{December 13, 2010}
\address{
Department of Mathematical Sciences\\
University of Copenhagen\\
Universitetsparken 5\\
DK-2100 Copenhagen E\\
Denmark}
\email{otogo@math.ku.dk}
\urladdr{http://www.math.ku.dk/~otogo}
\keywords{$K$-theory, $\KK$-theory}
\subjclass[2010]{Primary (19K35); Secondary (46L80)}
\maketitle

\begin{abstract} We extend McClure's results on the restriction maps in equivariant $K$-theory to bivariant $K$-theory:

Let $G$ be a compact Lie group and $A$ and $B$ be $G$-\cast-algebras. Suppose that $\KK^{H}_{n}(A, B)$ is a finitely generated $R(G)$-module for every $H \le G$ closed and $n \in \Z$. Then, if $\KK^{F}_{*}(A, B) = 0$ for all $F \le G$ {\em finite cyclic}, then $\KK^{G}_{*}(A, B) = 0$.
\end{abstract}

\setcounter{section}{-1}
\section{Introduction}

Let $G$ be a compact Lie group and let $X$ be a finite $G$-CW-complex. For any closed subgroup $H \le G$, we have a restriction functor in equivariant $K$-theory: 
	\begin{equation}
	\res^{G}_{H}: K_{G}(X) \to K_{H}(X).		
	\end{equation}

As an application of the generalized Atiyah-Segal completion theorem of \cite{MR935523}, McClure proved the following.
\begin{thm-i}[{McClure \cite[Theorem A]{MR862427}}] If $x \in K_{G}(X)$ restricts to zero in $K_{H}(X)$ for every {\em finite} subgroup $H$ of $G$, then $x = 0$. 
\end{thm-i}

Combining with Jackowski's result \cite[Corollary 4.3]{MR0448377}, one obtains the following.
\begin{thm-i}[{Jackowski-McClure \cite[Corollary C]{MR862427}}] 
If $K_{F}^{*}(X) = 0$ for all $F \le G$ finite cyclic, then $K_{G}^{*}(X) = 0$.
\end{thm-i}

We extend these to bivariant $K$-theory as follows. 

\begin{thm}\label{thm KK} Let $G$ be a compact Lie group and $A$ and $B$ be $G$-\cast-algebras. Suppose that $\KK^{H}_{n}(A, B)$ is a finitely generated $R(G)$-module for every $H \le G$ closed and $n \in \Z$.
\begin{enumerate}
\item[\thmC] 
Then, if $\KK^{F}_{*}(A, B) = 0$ for all $F \le G$ {\em finite cyclic}, then $\KK^{G}_{*}(A, B) = 0$.
\item[\thmA] Suppose, in addition, that $\KK^{F}_{n}(A, B)$ is a finitely generated group for all $F \le G$ finite and $n \in \Z$. Then, if $x \in \KK^{G}(A, B)$ restricts to zero in $\KK^{H}(A, B)$ for all $H \le G$ {\em finite}, then $x = 0$. 
\end{enumerate}
\end{thm}

\begin{rem}
\begin{enumerate}
\item See \cite{MR2083579} for a dual result for restriction maps in $K$-homology of spaces with actions of discrete groups.
\item Theorem \ref{thm KK} is in stark contrast to the results of Heath Emerson, where he showed that even for circle actions, noncommutative algebras can behave very differently from commutative ones. \cite{2010arXiv1004.2970E}
\end{enumerate}
\end{rem}

In fact, we prove the following. This is done mainly for clarity, but as an added bonus, we see that Theorem~\ref{thm KK} holds for equivariant $E$-theory as well.
\begin{thm}\label{thm L}
Let $G$ be a compact Lie group and let $\L_{G}^{*}$ be an $RO(G)$-gradable module theory over $\K_{G}^{*}$. Suppose that $\L_{H}^{n}(S^{0})$ is a finitely generated $R(G)$-module for every $H \le G$ closed and $n \in \Z$. Let $X$ be a finite based $G$-CW-complex.

\begin{enumerate}
\item[\thmC] Then, if $\L_{F}^{*}(X) = 0$ for all $F \le G$ {\em finite cyclic}, then $\L_{G}^{*}(X) = 0$.

\item[\thmA] Suppose, in addition, that $\L_{F}^{n}(S^{0})$ is a finitely generated group for all $F \le G$ finite and $n \in \Z$. Then, if $x \in \L_{G}^{*}(X)$ restricts to zero in $\L_{H}^{*}(X)$ for all  $H \le G$ {\em finite}, then $x = 0$. 
\end{enumerate} 
\end{thm}

The proof follows \cite{MR862427} very closely. In Section \ref{sec RO(G)}, we show that Theorem~\ref{thm L} implies Theorem~\ref{thm KK}. In Section \ref{sec Completion}, we extend the generalized Atiyah-Segal completion theorem of \cite{MR935523}, supplying the missing ingredient needed to finish the proof in Section \ref{sec Proof of main thm}. 

\begin{acknowledgments*} The author thanks the Centre for Symmetry and Deformation at the University of Copenhagen and the Danish National Research Foundation for support.
\end{acknowledgments*}

\section{$RO(G)$-graded cohomology theories}\label{sec RO(G)}

Let $G$ be a compact Lie group. A based $G$-space is a $G$-space with a $G$-fixed base point. In the rest of the paper, we assume that all $G$-spaces are $G$-CW-complexes and all cohomology theories are equivariant and reduced cohomology theories.

For a finite-dimensional representation $V$ of $G$, we write $S^{V}$ for the one-point compactification of $V$, considered a based $G$-space with base point the point at infinity.

\subsection{$RO(G;U)$-gradable theories}
We fix a complete universe $U$. (cf.\ \cite[Definition IX.2.1]{MR1413302}). 

\begin{defn} An {\em $RO(G)$-graded} cohomology theory is an $RO(G; U)$-graded cohomology theory in the sense of \cite[Definition XIII.1.1]{MR1413302}. A {\em $\Z$-graded} cohomology theory is an $RO(G; U^{G})$-graded cohomology theory (any trivial universe would work). We say that a $\Z$-graded cohomology theory is {\em $RO(G)$-gradable} if it is the $\Z$-graded part of an $RO(G)$-graded theory.
\end{defn}


Let $\E_{G}^{*}$ be a $\Z$-graded cohomology theory. 
For a closed subgroup $H \le G$ and a based $H$-CW-complex $X$, we define 
	\begin{equation}
	\E_{H}^{*}(X) \coloneqq \E_{G}^{*}(G_{+} \wedge_{H}X).	
	\end{equation}
Then $\E_{H}^{*}$ is a $\Z$-graded cohomology theory on based $H$-spaces. If $X$ is actually a based $G$-CW-complex, then we have a natural $G$-equivariant  identification 
	\begin{equation}
	G_{+} \wedge_{H}X \cong G/H_{+} \wedge X
	\end{equation}
and the collapse map $G/H \to \pt$ gives rise to a natural transformation
	\begin{equation}
	\res^{G}_{H}: \E_{G}^{*} \to \E_{H}^{*}
	\end{equation}
called the {\em restriction} map. 

\subsection{Bivariant $K$-theory} The following is the main example we have in mind. First note that $\K_{G}^{*}$ is an $RO(G)$-graded commutative ring theory with $\K_{G}^{V}(X) = \KK^{G}(C_{0}(S^{V}), C_{0}(X))$ and $R(G) \cong \K_{G}(S^{0})$.

\begin{prop} Let $G$ be a compact Lie group and let $A$ and $B$ be $G$-\cast-algebras. For a finite based $G$-CW-complex $X$ and finite-dimensional real representation $V$ of $G$, we define
	\begin{equation}
	\L_{G}^{V}(X) \coloneqq \KK^{G}(A \otimes C_{0}(S^{V}), B \otimes C_{0}(X)).
	\end{equation}
Then the following holds.
\begin{enumerate}[(i)]
\item $\L_{G}^{*}$ defines an $RO(G)$-graded cohomology theory on the category of finite based $G$-CW-complexes.
\item $\L_{G}^{*}$ extends to an $RO(G)$-graded cohomology theory on the category of based $G$-CW-complexes.
\item $\L_{G}^{*}$ is a module theory over $\K_{G}^{*}$.
\end{enumerate}
\end{prop}
\begin{proof}
(i) See \cite{MR918241}.
(ii) By Adams' representation theorem \cite[Theorem XIII.3.4]{MR1413302}, $\L_{G}^{*}$ is represented by an $\Omega$-$G$-prespectrum, hence extends to an $RO(G)$-graded cohomology theory on the category of $G$-CW-complexes. See \cite{MR1193150}.
(iii) The module structure
	\begin{equation}
	\L_{G}^{V}(X) \times \K^{W}_{G}(Y) \to \L_{G}^{V+W}(X \wedge Y).
	\end{equation}
is given by the Kasparov product 
	\begin{align}
	\KK^{G}(A(S^{V}), B(X)) \times \KK^{G}(C_{0}(S^{W}), C_{0}(Y))\\ \to \KK^{G}(A(S^{V+W}), B(X \wedge Y)).
	\end{align}
\end{proof}

It is well-known that for $H \le G$,  
	\begin{equation}
	\KK^{G}(A, B \otimes C_{0}(G/H_{+})) \cong \KK^{H}(A, B)	
	\end{equation}
and the restriction map is induced by $G/H_{+} \to S^{0}$. Hence we obtain the following corollary.

\begin{cor}
Suppose that Theorem~\ref{thm L} holds. Then Theorem~\ref{thm KK} holds.
\qed
\end{cor}
\section{Atiyah-Segal Completion}\label{sec Completion}

First we abstract the main finiteness condition from Theorem~\ref{thm L}.
\begin{defn} Let $R$ be a unital commutative ring and let $\L_{G}^{*}$ be a $\Z$-graded cohomology theory with values in $R$-modules. We say that $\L_{G}^{*}$ is {\em finite} over $R$ if $\L_{G}^{n}(X)$ is a finitely generated $R$-module for every {\em finite} based $G$-CW-complex $X$ and $n \in \Z$.
\end{defn}

Clearly, this is equivalent to asking that $\L_{H}^{k-n}(S^{0}) \cong \L_{G}^{k}(G/H_{+} \wedge S^{n})$ is a finitely generated $R$-module for $H \le G$.

\begin{lem}\label{lem cohomology} Let $G$ be a compact Lie group and let $R$ be a unital commutative ring. Let $\L_{G}^{*}$ be a $\Z$-graded cohomology theory with values in $R$-modules. Suppose that $R$ is Noetherian and $\L_{G}^{*}$ is finite over $R$. Then for any family $\I$ of ideals in $R$, the following defines a $\Z$-graded cohomology theory with values in pro-$R$-modules:
	\begin{equation}
	\L_{G}^{*}(X)^{\wedge}_{\I} \coloneqq \{\L_{G}^{*}(Y)/J \cdot \L_{G}^{*}(Y)\}.
	\end{equation}
where $Y \subseteq X$ runs over the finite based $G$-CW-subcomplexes of $X$ and $J$ runs over the finite products of ideals in $\I$.  
\end{lem}
Note that in this lemma, it is enough to have $\L_{G}^{*}$ to be a cohomology theory on {\em finite} based $G$-CW-complexes (only finite wedges are considered in the additivity axiom). 
\begin{proof} Exactness follows from the Artin-Rees lemma. See the proof of \cite[Lemma 2.1]{MR935524}.
\end{proof}

\subsection{Bott Periodicity}
Let $V$ be a complex $G$-representation. By Bott periodicity \cite[Theorem 4.3]{MR0228000}, $\K_{G}^{0}(S^{V})$ is a free $\K_{G}^{0}(S^{0})$-module generated by the Bott element $\lambda_{V} \in \K_{G}^{0}(S^{V})$. The {\em Euler class} of $V$ is defined to be $e^{*}(\lambda_{V}) \in \K_{G}^{0}(S^{0})$, where $e: S^{0} \to S^{V}$ is the obvious map. 

\begin{lem}\label{lem Bott} Let $\L_{G}^{*}$ be an $RO(G)$-graded module theory over $\K_{G}^{*}$. Then for any complex representation $V$, multiplication by the Bott element $\lambda_{V} \in \K_{G}^{0}(S^{V})$ gives an isomorphism 
	\begin{equation}
	\L_{G}^{0}(S^{0}) \cong \L_{G}^{0}(S^{V}). 
	\end{equation}
If $V \subseteq W$ are complex representations and $i: S^{V} \to S^{W}$ is the inclusion, then the following diagram commutes
	\begin{equation}
	\xymatrix{
	\L_{G}^{0}(S^{0}) \ar[r] \ar[d]^{\cdot \chi_{W-V}}& \L_{G}^{0}(S^{W}) \ar[d]^{i^{*}}\\
	\L_{G}^{0}(S^{0}) \ar[r] & \L_{G}^{0}(S^{V})\\
	}.
	\end{equation}
\end{lem}
\begin{proof} Let $\lambda_{V}^{-1} \in \K_{G}^{V}(S^{0})$ denote the inverse Bott element: it has the property that
	\begin{equation}
	\lambda_{V} \cdot \lambda_{V}^{-1} = \lambda_{V}^{-1} \cdot \lambda_{V} = 1 \in \K_{G}^{V}(S^{V}) \cong \K_{G}^{0}(S^{0}).
	\end{equation}
Then multiplication by $\lambda_{V}^{-1}$ gives the inverse map
	\begin{equation}
	\L_{G}^{0}(S^{V}) \to \L_{G}^{V}(S^{V}) \cong \L_{G}^{0}(S^{0}).
	\end{equation}
The second statement is shown for $\L_{G}^{*} = \K_{G}^{*}$ in \cite[page 4]{MR935523}. The general case follows by functoriality.
\end{proof}
	
\subsection{Completion}
A class of subgroups of $G$ closed under subconjugacy is called a {\em family}. A family $\C$ of subgroups of $G$ determines a class, again denoted  $\C$, of ideals of $R(G)$ by the kernels of the restriction maps:
	\begin{equation}
	\ker(\res^{G}_{H}: R(G) \to R(H)),\quad H \in \C,	
	\end{equation}
hence a topology on any $R(G)$-module. 

The following is a straightforward generalization of \cite[Theorem 3.1]{MR935523}.
\begin{thm}\label{thm AS} Let $G$ be a compact Lie group and let $\L_{G}^{*}$ be an $RO(G)$-gradable module theory over $\K_{G}^{*}$, which is finite over $R(G)$. 

Let $\C$ be a family of subgroups of $G$. For any based $G$-CW-complex $X$, if $\L_{H}^{*}(X)^{\wedge}_{\C|H} = 0$ for all $H \in \C$, then $\L_{G}^{*}(X)^{\wedge}_{\C} =0.$

\end{thm}
\begin{proof} By \cite[Corollary 3.3]{MR0248277}, $R(G) = \K_{G}^{0}(S^{0})$ is Noetherian. Hence, by Lemma~\ref{lem cohomology}, $\L_{G}^{*}(X)^{\wedge}_{\C}$ is a cohomology theory.

Now the proof of \cite[Theorem 3.1]{MR935523} carries over ad verbatum, once we extend Bott periodicity to $\L_{G}^{*}$ as in  Lemma~\ref{lem Bott}.
\end{proof}

\begin{cor}\label{cor AS}
Let $E\C$ denote the classifying space of $\C$. For any finite based $G$-CW-complex $X$, the projection map $E\C_{+} \to S^{0}$ gives completion
	\begin{equation}
	\L_{G}^{*}(E\C_{+} \wedge X) \cong \lim \L_{G}^{*}(Y \wedge X) \cong \lim \L_{G}^{*}(X)^{\wedge}_{\C},
	\end{equation}
where $Y$ runs over finite based subcomplexes of $E\C_{+}$. 	
\end{cor}
\begin{proof}
The inverse system $\L_{G}^{*}(X)^{\wedge}_{\C}$ satisfies the Mittag-Leffler condition and $\L_{G}^{*}(Y \wedge X)$ is $\C$-complete for any finite based subcomplex $Y \subset E\C_{+}$ (cf.\ \cite[Corollary 2.1]{MR935523}).
\end{proof}

\section{Proof of Theorem \ref{thm L}}\label{sec Proof of main thm}
\subsection{$\F$-spaces}

Let $\F$ be a family of subgroups of $G$. We say that a based $G$-CW-complex $X$ is an $\F$-space if all the isotropy groups, except at the base point, are in $\F$. The following lemma says that in the proof of Theorem~\ref{thm L}, we may assume that $X$ is an $\F$-space, for any $\F$ containing all finite cyclic subgroups of $G$.

\begin{lem}\label{lem F-space} 
Let $G$ be a compact Lie group and let $\L_{G}^{*}$ be an $RO(G)$-gradable module theory over $\K_{G}^{*}$, which is finite over $R(G)$. 

Let $\F$ be a family containing all finite cyclic subgroups of $G$. Then for any finite based $G$-CW-complex $X$,  the top horizontal map in the commutative diagram
	\begin{equation}
	\xymatrix{\L_{G}^{*}(X) \ar[r] \ar[d] & \lim_{Y \subset E\F_{+}}\L_{G}^{*}(Y \wedge X) \ar[d]\\
	\prod_{F \in \F} \L_{F}^{*}(X) \ar[r] &\lim_{Y \subset E\F_{+}}\prod_{F \in \F} \L_{F}^{*}(Y \wedge X)\\ 
	},
	\end{equation}
is injective. Here $Y$ runs over the finite based subcomplexes of $E\F_{+}$, the horizontal maps are induced by the projections $Y \wedge X \to X$ and the vertical maps are restrictions. 
\end{lem}
\begin{proof} The $\F$-topology on $\L_{G}^{*}(X)$ is Hausdorff by \cite[Corollary 3.3]{MR862427}. Hence, the claim follows from Corollary~\ref{cor AS}.
\end{proof}

Let $\C$ denote the family of finite cyclic subgroups of $G$.

\begin{proof}[Proof of Theorem~\ref{thm L}\thmC]
By assumption, $\L_{F}^{*}(X) = 0$ for all $F \in \C$. Let $Y$ be a finite based $G$-CW-complex, which is a $\C$-space. Then the zero skeleton $Y^{0}$ and the skeletal quotients $Y^{n}/Y^{n-1}$ are finite wedges of $G$-spaces of the form $G/F_{+} \wedge S^{n}$ with $F \in \C$. It follows that $\L_{G}^{*}(Y \wedge X) = 0$. Hence by Lemma~\ref{lem F-space}, $\L_{G}^{*}(X) = 0$.
\end{proof}

\subsection{Induction}
We write $\O_{G}$ for the category whose objects are orbit spaces $G/H$, where $H \le G$ is a closed subgroup, and whose morphisms are homotopy classes of $G$-maps. 

Recall that a compact Lie group is said to {\em cyclic} if it has a topological generator (an element whose powers are dense) and {\em hyperelementary} if it is an extension of a cyclic group by a finite $p$-group.

We write $\H$ for the class of hyperelementary subgroups of $G$ and let $\O_{\H}$ denote the full subcategory of $\O_{G}$ of orbits $G/H$ with $H$ subconjugate to a subgroup in $\H$.

\begin{lem}\label{lem ind} Let $G$ be a compact Lie group and let $\L_{G}^{*}$ be an $RO(G)$-gradable module theory over $\K_{G}^{*}$. Then, for any based $G$-CW-complex,  the restriction maps induce an isomorphism
	\begin{equation}
	\L_{G}^{*}(X) \cong \lim_{\O_{\H}} \L_{H}^{*}(X).
	\end{equation}
\end{lem}
\begin{proof} Follows from Propositions 2.1 and 2.2 of \cite{MR862427}.
\end{proof}

For any abelian group $M$, let $M^{\wedge}_{\Z}$ denote its adic completion $\lim_{n} M/nM$.

\begin{proof}[Proof of Theorem \ref{thm L}\thmA]
Let $\F$ denote the family of finite subgroups of $G$.

By Lemma \ref{lem ind}, we may assume that $G$ is a hyperelementary group and by Lemma \ref{lem F-space}, we may assume that $X$ an $\F$-space. 

Let $G$ be a hyperelementary group and $X$ an $\F$-space. Then the restriction map 
	\begin{equation}
	\L_{G}^{*}(X)^{\wedge}_{\Z} \to \lim_{F \in \O_{\F}} \L_{F}^{*}(X)^{\wedge}_{\Z}.
	\end{equation}
is an isomorphism by \cite[Theorem 1.1]{MR862427}. By \cite[Corollary 3.3]{MR862427}, the adic topologies on $\L_{G}^{*}(X)$ and $\L_{F}^{*}(X)$ are Hausdorff. This completes the proof.
\end{proof}
\bibliographystyle{amsalpha}
\bibliography{../BibTeX/biblio}
\end{document}